OpenStudy (anonymous):

Part 1 Choose the graph of the linear system. x – y = 1 x + y = –5

OpenStudy (anonymous):

Part 2 What is the solution of the linear system?

OpenStudy (uri):

x – y = 1 x + y = –5 ========= x=-4 -y and +y will be cancelled.

OpenStudy (uri):

Now you have to find y Put the value of x in the 2nd eq x + y = –5 -4+y-5 y=-5+4 y=-1 (-4,-1)

OpenStudy (anonymous):

Options

OpenStudy (adilalvi):

a

OpenStudy (anonymous):

Thanks Uri and Adilalvi

OpenStudy (uri):

Yw :)

OpenStudy (anonymous):

So it would be option 1 right?

OpenStudy (abb0t):

you know, the equation of a line is defined as: y=mx+b you can always rearrange it in that form to see the line... where "b" is your "y-coordinate" m is your slope, which is basically \(\frac{rise}{run}\)

OpenStudy (adilalvi):

yes exactly